OpenStudy (sleepyjess):

Can someone check this mathematical induction problem? Question and work below

OpenStudy (sleepyjess):

1^`2 + 4^2 + 7^2 + ... + (3n - 2)^2 = \(\dfrac{n(6n^2-3n-1)}2\)

OpenStudy (sleepyjess):

1^2 = \(\dfrac{1(6(1)^2 - 3(1)-1)}2\)

OpenStudy (sleepyjess):

\(1^2 = \dfrac{1(36-3-1)}2\)

OpenStudy (sleepyjess):

\(1^2 = \dfrac{36}2\)

OpenStudy (sleepyjess):

\(1 \ne 16\)

OpenStudy (alexandervonhumboldt2):

mistake: trird post: 1*(6*(1)^2-3*(1)-1)=1*(6*1-6-1)

OpenStudy (alexandervonhumboldt2):

you got as 6*(1)^2=36 but that not true

OpenStudy (sleepyjess):

oh whoops, thanks for catching that :)

OpenStudy (alexandervonhumboldt2):

np

OpenStudy (sleepyjess):

So really it would be 2/2 = 1 which is true

OpenStudy (alexandervonhumboldt2):

yes that correct and true

OpenStudy (sleepyjess):

Thank you ^_^

OpenStudy (alexandervonhumboldt2):

np